OpenStudy (anonymous):

I'll FAN & MEDAL!! What is the sum of the first eight terms of a geometric series whose first term is 3 and whose common ratio is 1/2 ?

OpenStudy (solomonzelman):

my connection has been very bad lately. finally got to the thread;)

OpenStudy (solomonzelman):

For a geometric sequence, starting from the first term \(a1\) and a common ratio \(\rm r\), the sum for \(\rm n\) number of terms (in the sequence) is given by the following formula. \(\large\color{black}{ \displaystyle {\rm S}_{\rm n}=\frac{a_1\times \left(1-{\rm r}^{\rm n}\right) }{1-{\rm r}} }\)

OpenStudy (solomonzelman):

You need to add 8 terms, so your n is 8. ----------------------------------- Your common ratio is ½, so r=½. ----------------------------------- And your first term is 3, so \(a_1\)=3 -----------------------------------

OpenStudy (solomonzelman):

Just plug this information into the formula, and calculate. Good luck:)